OpenStudy (anonymous):

I do not understand this at all! Help please?? What is the 13th term in the Fibonacci Sequence? Term 1 2 3 4 5 6 7 8 9 Fibonacci Number 1 1 2 3 5 8 13 21 34

jimthompson5910 (jim_thompson5910):

To get the 10th term, add the last two terms (terms #8 and #9) tenth term: 21+34 = 55 To get the 11th term, repeat, but do it for terms #9 and #10 eleventh term: 34+55 = 89 Keep going until you hit the 13th term

OpenStudy (anonymous):

so 55+89= 144 is the 12th 89 +144= 233??

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

what does it mean by whats the ration of 13th and 14th?

jimthompson5910 (jim_thompson5910):

find the 14th term then divide the 14th term by the 13th term

OpenStudy (anonymous):

377/144= 2.61805? it says 5 decimal places so would that be right?

jimthompson5910 (jim_thompson5910):

14th term: 377 13th term: 233 (NOT 144) Ratio of 14th term and 13th term: (14th term)/(13th term) = 377/233 = 1.61803

jimthompson5910 (jim_thompson5910):

Side note (probably optional), the number 1.61803 is a special number and it is known as the golden ratio (it's like pi where it goes on forever)

OpenStudy (anonymous):

Ohh! thank you soo muuch!!!

jimthompson5910 (jim_thompson5910):

np